### v1

# Exploring Parent’s Needs for Children-Centered AI to Support Preschoolers’ Storytelling and Reading Activities

Abstract.
Interactive storytelling is vital for preschooler development. While children’s interactive partners have traditionally been their parents and teachers, recent advances in artificial intelligence (AI) have sparked a surge of AI-based storytelling technologies. As these technologies become increasingly ubiquitous in preschoolers’ lives, questions arise regarding how they function in practical storytelling scenarios and, in particular, how parents, the most critical stakeholders, experience and perceive these technologies. This paper investigates these questions through a qualitative study with 17 parents of children aged 3-6. Our findings suggest that even though AI-based storytelling technologies provide more immersive and engaging interaction, they still cannot meet parents’ expectations due to a series of interactive, functional, and algorithmic challenges. We elaborate on these challenges and discuss the possible implications of future AI-based storytelling technologies for preschoolers. We conclude by highlighting the design implications for future AI-based storytelling technologies.
Storybook reading, storytelling, interactive, AI, preschoolers, artificial intelligence, parents
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June XX–$X$, 20XX; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCI
1. Introduction
Interactive storytelling together with the parents is a critical step in the development for preschoolers111”Preschoolers” in this paper refers to young children, typically aged 3 to 6, who are in the stage of early childhood education that precedes formal schooling (of Disease Control and Prevention, 2021). (Glassner, 2017). Because preschoolers are often not able to decode text independently, parents often sit together with and read to them, ideally engaging the children in guided conversation (Zevenbergen and Whitehurst, 2003) by asking questions and providing responsive feedback. This process has been proven to be beneficial to children’s later language and literacy proficiency, communication skills, cultural and emotional awareness, and other aspects of cognitive development (Wright, 1995; Yen et al., 2018; Chang and Huang, 2016; Peck, 1989). However, due to various practical challenges, today’s parents may not always have sufficient language skills, knowledge, time, or inclination to engage in such conversation-rich storybook reading with their preschoolers, even though they believe in the importance of interactive storytelling with their children (Manz et al., 2010; Xu et al., 2021b).
In recent years, AI-based applications with multiple functions, materials, modalities, and interaction techniques have been becoming increasingly popular in supporting storytelling practices for preschoolers in the household setting. These applications could be embedded as applications in mobile phones or tablets (Dietz et al., 2021; Zhang et al., 2022), or exist with physical tangible presences (e.g., Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021)); they rely on interactive technologies tailored to children (e.g., (Xu et al., 2022; Yao et al., 2022)), or general voice assistant technologies (e.g., Amazon Alexa, Google Assistant, Apple Siri); they provide specialized features designed specifically for children’s storytelling (e.g., (Dietz et al., 2021; Zhang et al., 2022; Michaelis and Mutlu, 2017)), or carry out children’s instructions based on the embedded general voice assistant technologies (e.g., Amazon Alexa, Google Assistant, Apple Siri); some of these applications are mature commercial products (e.g., Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021)) and some others are still in the research period (e.g., storycoder (Dietz et al., 2021), and Storybuddy (Dietz et al., 2021; Zhang et al., 2022)).
With increasingly powerful and intelligent capabilities, these AI-based systems and tools have provided tremendous promises for interactive, intelligent, and engaging storytelling experiences for children, acting as their language partners or learning assistants (Xu et al., 2021b; Dietz et al., 2021). Recent booming large-language model-based technologies (LLMs) (Radford et al., 2019) further solidify this vision. This increasing functionality, diversity, and popularity of AI-based technologies in storytelling scenarios for preschoolers raise critical questions regarding: how AI techniques support real-world storytelling practices and how parents, the vital role of storytelling practices, experience, and perceive AI-mediated storytelling for preschoolers?
In CSCW, studies have paid attention to the design and use of various storytelling applications, with the aim of improving children’s digital literacy (Dietz et al., 2021; Tseng et al., 2021), stimulating their creativity and collaboration (Dietz et al., 2021; Helmes et al., 2009), fostering their connection and interaction with parents (Harbig et al., 2011; Zhang et al., 2022), etc. Meanwhile, researchers have evaluated the functionality and effectiveness of existing commercial AI products (e.g., Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021)) as well as the recent LLMs technologies (Han and Cai, 2023b) to support storytelling practices with preschoolers (Xu et al., 2021b; Garg et al., 2022; Chubb et al., 2022). However, these studies were conducted primarily through experimental research. To date, very little has been published on the usage and evaluation of AI-based storytelling technologies in real-world storytelling scenarios beyond a limited research period.
More importantly, while these studies have paid attention to accessing the effectiveness and usability of AI technologies in children’s growth and development, the experience, perceptions, and expectations of parents, who actually play the most significant roles in practical storytelling (Frude and Killick, 2011; Vasalou et al., 2020) with the children, have been largely overlooked. In particular, given the cognitive and social development stage, preschoolers often lack the essential AI literacy (Su et al., 2023) and skills (Lever and Sénéchal, 2011; Resnick et al., 2009) to independently use various AI technologies, thus relying on parents to navigate this process (Zhang et al., 2022). Neglecting the understanding of parents’ needs and perceptions may significantly hinder the effectiveness of these technologies in real-world contexts.
Our study fills this research gap through a qualitative study of 17 parents with preschoolers between the ages of 3 and 6. Our findings indicate that parents have a cautiously positive attitude toward AI-based storytelling. Given the practical challenges facing today’s parents, such as lack of time, energy, and necessary knowledge in carrying out high-quality, meaningful interactive storytelling, our findings suggest that parents expect AI technologies to help alleviate their burdens and support the desired parent-children interactive storytelling. Yet, existing tools generally fail to support parents’ expected storytelling in the real world due to several interactive and algorithmic challenges. Drawing on our findings, we delve into parents’ expectations for ideal parent-child interactive storytelling and propose several crucial design implications for future research.
The key contributions of our work are mainly twofold:
  • We acquire an empirical understanding of parents’ perception and facing challenges regarding the use of AI-based storytelling tools in practice through a qualitative study involving 17 parents with preschoolers.
  • We identify and analyze vital design implications and recommendations to better facilitate the increasing demand for AI-based children-centered storytelling.
2. Related Work
Our study situates into the research of AI-based storytelling for preschoolers and its practical effectiveness. In this section, we first present a review of existing AI-based storytelling tools. Following that, we review existing literature about the practical usage and evaluation of these tools from children’s and parents’ perspectives respectively.
2.1. AI-Based Interactive Storytelling Systems
With increasingly powerful and intelligent capabilities, as well as the practical barriers and challenges in traditional parent-child storytelling, there has been a growing CSCW and HCI interest in designing and using AI-based technologies in children’s storytelling scenarios (Frohlich et al., 2009; Dietz et al., 2021; Tseng et al., 2021; Helmes et al., 2009; Harbig et al., 2011; Zhang et al., 2022; Rubegni and Landoni, 2014). Researchers, for instance, have designed storyCoder (Dietz et al., 2021), a voice-guided smartphone application, and PlushPal (Tseng et al., 2021), a web-based tool, to improve children’s computational thinking ability and data literacy by leveraging storytelling as a creative activity. Helmes et al. (2009) designed TellTable, a tabletop storytelling application, to stimulate children’s creativity and collaboration. Meanwhile, researchers have paid specific attention to children with disabilities, such as Gelsomini et al. (2016)’s Wildcard for children with intellectual developmental disabilities and Harbig et al. (2011)’s SignBright for deaf children. Besides, commercially available AI-based robotic toys, such as Alpha Egg (Toycloud, [n. d.]), Luka (Zhao and McEwen, 2021), Codi (Learning, 2021), etc., also provide platforms for facilitating interactive storytelling experiences for children (Garg and Sengupta, 2020).
The recent booming of Large Language Model technologies (LLMs) further sparked the potential of AI-based storytelling technologies in supporting more intelligent and personalized storytelling practices. Han and Cai (2023b), for instance, examines the design implications of leveraging generative AI tools for improving children’s literacy development and creative story expression. Based on the findings, they proposed AIStory (HAN et al., 2023), a generative AI-powered visual storytelling application prototype that can be used for children’s creative expression, storytelling, and literacy development. Li and Xu (2023) propose a peer-like embodied conversational agent named STARie that integrates multiple AI models that aims to support narrative production in collaborative storytelling, specifically for children aged 4-8.
While these studies provide the worth-expecting vision for AI-mediated storytelling, most existing technologies are still in the prototype design and evaluation phase, involving brief testing for a short time with limited users rather than applying them to real-world scenarios for a long time. Due to the lack of large-scale deployment and long-term usage, it remains unclear how such AI-based storytelling systems are perceived and used by their intended users (e.g. parents and preschoolers) and what barriers and challenges they might encounter. Our study contributes to bridging this gap by examining the use of existing AI-based storytelling tools in the real world, particularly from parents’ perspective.
2.2. AI-Based Storytelling Technologies and Children
In CSCW and related fields, there has been surging interest in evaluating the performance of AI-based intelligent systems in children’s reading and storytelling scenarios. Researchers, for instance, have illustrated that preschool-aged children can generally interact naturally with conversational agents (CAs), perceiving them as both intelligent and friendly (Druga et al., 2017). Interacting with CAs can also improve children’s story comprehension and learning, comparable to that provided by an adult language partner.(Xu et al., 2021b; Kory and Breazeal, 2014; Beneteau et al., 2020; Xu et al., 2021a; Xu and Warschauer, 2020a). Compared to reading and interacting with human partners, children reading with CAs responded to questions with better intelligibility but lower productivity, lexical diversity, and topical relevance (Xu et al., 2021b).
Meanwhile, literature has also shown that AI applications face difficulties when integrating into real-world storytelling or education practices for preschoolers (Su et al., 2023). From the technical standpoint, researchers, for instance, have demonstrated existing conversational agents are hard to generating the expected valid and reliable comprehension story questions and supporting tailored storytelling and interaction (Xu et al., 2022; Yao et al., 2022; Bentley et al., 2018; Cowan et al., 2017; Luger and Sellen, 2016), which caused children might encounter difficulties in understanding questions from agents, fail to follow the conversation flow, and struggle with appropriate turn-taking when interacting with agents (Du et al., 2021; Pantoja et al., 2019). In addition, many existing AI-based storytelling tools merely ask questions in a sequential manner from a pre-generated list, which does not prioritize child engagement, as it lacks logical coherence between questions and often fails to provide guidance to children in case of confusion or incorrect answers (Zhang et al., 2022; Xu et al., 2022; Yao et al., 2022).
As children are increasingly exposed to various AI technologies, researchers have also explored barriers and limitations of AI technologies to children from a broader perspective of child-AI interaction. They, for instance, examined the potential social and gender biases of AI technologies to children, and proposed to provide them with opportunities and skills to critically reflect on current AI applications and envision (Sharma et al., 2023; Zhou and Yu, 2023). Meanwhile, researchers have also examined the impact of increasing digital technologies and the trend of datafication on children (Wang et al., 2022a, 2023). They suggested future AI technologies should be more transparent and autonomy-supportive to children. Given these children-centered findings, researchers proposed to adapt a generic human-centered AI design framework (Wang et al., 2022b; Zhou and Yu, 2023) for more ethical, inclusive, and fair AI technology designs. These findings and claims could also be reflected in AI storytelling technologies for preschoolers.
Overall, the challenges and limitations identified in prior research raise an urgent need to thoroughly understand users’ usage and experience of AI-based storytelling technologies, and identify the practical challenges and barriers. In particular, preschoolers often lack the essential AI literacy (Su et al., 2023) and skills (Resnick et al., 2009) to independently use AI technologies. These insights are, therefore, more critical in informing the development of future AI-based storytelling technologies for preschoolers, ensuring the alignment with practical storytelling needs, and maximizing benefits in AI-based storytelling. Our work joins the effort of ongoing work to deeply explore the practical challenges and limitations of AI-based storytelling technologies.
2.3. AI-Based Storytelling Technologies and Parents
For practical storytelling practices with preschoolers, parents play the most vital role in, for instance, shaping children’s language and comprehension skills development, and strengthening the bond between parents and children (Frude and Killick, 2011; Vasalou et al., 2020). In AI-based storytelling practices for preschoolers, parents’ roles and functions are even more significant, because preschoolers often lack the essential literacy (Su et al., 2023) and skills (Resnick et al., 2009) to independently use various AI technologies, and they therefore often rely on parents to navigate this AI-based storytelling process (Zhang et al., 2022). According to the study of Lin et al. (2021), when integrating AI-based intelligent agents into the story reading and telling scenario, about 63% interactions are actually parent-driven, with the majority of parents guiding children by redirecting their attention or clarifying virtual agents’ prompts.
However, researchers have found that existing AI-based storytelling technologies often excluded the involvement of parents (Zhang et al., 2022), and the evaluation of these technologies has also mainly from the children’s needs and perspective, paying attention to children’s cognition of AI technologies and the impact of AI technologies on children’s certain abilities, e.g., (Xu et al., 2021b; Dietz et al., 2021; Cai, 2023a; Gelsomini et al., 2016; Chubb et al., 2022). In general, attention to parents’ experiences and perceptions of AI-based storytelling is limited. In CSCW and HCI, some prior research has provided formative insights into parents’ objectives for interactive storytelling. Zhang et al. (2022), for instance, find that parents appreciate the educational potential of new technologies in enhancing parent-child interaction and actively seeking customized content for their children. Boffi (2020) reports the challenges parent face in educating their children to interact with intelligent storytelling systems correctly and respectfully. Apart from these studies, little prior work has deeply explored parents’ experiences and perceptions of AI-based storytelling technologies.
Given parents’ significant roles in the adopting and using of AI-based storytelling systems (Boffi, 2020; Lin et al., 2021; Zhang et al., 2022), our study provided an empirical study to comprehensively and deeply examine how they perceive and use existing AI-based storytelling tools, their encountered barriers and challenges. Relevant to our study here is Lin et al. (2021)’s study about parental acceptance of children’s storytelling robots. Through a qualitative study with 18 parents, they demonstrate parents’ mixed, though generally positive, attitudes toward children’s storytelling robots. In contrast to their research, our study primarily focused on the challenges and challenges our participants experienced and perceived in real-world settings. Compared to experimental research which occurs in controlled environments, and design fiction (Lindley and Coulton, 2015) which uses fictional scenarios and prototypes to explore and envision possible futures for technology, studying the lived experiences of parents in real-world scenarios can provide us valuable insights into parents’ actual use and perceptions of AI storytelling tools, allowing us to capture the nuances and complexities of how parents integrate such tools into their daily storytelling scenarios, and uncover unforeseen challenges and benefits.
In addition, we paid our attention to more general AI-based storytelling technologies, instead of the specific storytelling robots as what Lin et al. (2021) have done. During our study, we did not specify what kinds of technologies our participants have used, so that the questions could be inclusive to participants to share various levels of experiences with different technologies mediated storytelling practices. Through this broader research scope, we hope to contribute to the general AI-based storytelling technology design by informing designers and developers with in-depth and empirical understandings of the practical challenges and challenges in fitting these technologies into the practical parent-child storytelling scenario and helping them better address these challenges.
3. METHOD
This paper aims to gain insights into parents’ practical adoption, usage, experiences, and perceptions toward AI-based intelligent storytelling technologies and tools for preschoolers between the ages of 3 and 6. We are particularly interested in investigating 1) current acceptance and usage of AI-based storytelling technologies and tools by parents with preschoolers, 2) their perceived benefits, 3) encountered challenges in using these tools, and 4) their perceptions and expectations of future AI-based intelligent storytelling technologies. To this end, we conducted a qualitative interview study with 17 parents. The rest of this section details the participants and recruitment, the types of data being collected, and the data analysis process.
3.1. Participants and Recruitment
We looked for participants who 1) were parents with preschoolers between the ages of 3 and 6 and 2) had previous or ongoing experience using AI-based intelligent tools to support storytelling activities. We did not specify what AI tools should be in our opinion and what kinds of AI tools participants had used. We recruited ones as long as they considered they had used storytelling tools with AI features for their storytelling practices with preschoolers. We also did not specify how much experience they should have, so that the recruitment could be open to people with various levels of usage and engagement in various kinds of AI-based storytelling technologies.
The participants were recruited via a snowball sampling approach. The first author, a native Mandarin speaker, contacted her classmates and friends who had preschoolers. After confirming that s/he had used at least one kind of AI-based storytelling tool, s/he then disclosed our research intention and recruited her/him for our interviews. Seven were contacted, and all agreed to be interviewed. To avoid bias from our own social networks, we then sought more participants by posting recruitment messages to online forums and communities. We designed the recruitment poster including self-introduction, research intention, recruitment criteria, and contact information of the first author, and team members distributed the poster to their social networks. Twelve reached out directly, and 10 of them agreed to be interviewed.
ID Roles Children Education Work Used Tools
P1
Mother
4-yo son, 10-yo daughter
Finance, Econ.
Company Staff
Interactive Talking Pen (C1), Robot (C5), Little Genius (C3), XiaoDu (C6)
[…]
[0.5pt/0.5pt] P2
Mother
6-yo daughter, 12-yo son
Comp. Sci.
High School Staff
HuoHuoTu (C2), iPad (C4), Xiaodu (C6), Tmall Genie (C6)
[…]
[0.5pt/0.5pt] P3
Father
5-yo daughter
Health Care
Company Employee
Talking Pen (C1), HuoHuoTu (C2), Early Learning Device (C3), iPad (C4), XiaoDu (C6)
[…]
[0.5pt/0.5pt] P4
Father
5-yo daughter
Business Admin.
Company Manager
Jiaojiao (iPad app) (C4), Qiaohu (an educational brand, including TV programs, animated series, interactive games, and educational materials)
[…]
[0.5pt/0.5pt] P5
Mother
6-yo daughter
Automation
CIO
Interactive Pen (C1), Interactive Player (C2), iPad (C4), Robot (C5)
[…]
[0.5pt/0.5pt] P6
Mother
6-yo daughter
Business Admin.
Data Company Employee
HuoHuoTu (C2), Smartwatch with story reading apps, iPad (C4)
[…]
[0.5pt/0.5pt] P7
Mother
5-yo son
Educational Psychology
University Staff
Robot (C5)
[…]
[0.5pt/0.5pt] P8
Mother
6-yo son
Comp. Sci.
Researcher
Interactive Talking Pen (C1), Seewo (C3), Luka (C5), iPad (C4)
[…]
[0.5pt/0.5pt] P9
Mother
6-yo son
Finance, Econ.
University Financial Staff
Interactive Talking Pen (C1)
[…]
[0.5pt/0.5pt] P10
Mother
5-yo daughter
Public Admin.
Company Employee
Qiaohu
[…]
[0.5pt/0.5pt] P11
Father
4-yo daughter
Comp. Sci.
University Researcher
Interactive Pen (C1), Early Learning Device (C3), Robot (C5), Siri (C6), Qiaohu
[…]
[0.5pt/0.5pt] P12
Father
6-yo daughter
Education
University Staff
Tmall Genie (C6), XiaoAi (C6), iPad (C4)
[…]
[0.5pt/0.5pt] P13
Father
3-yo son, 5-yo daughter
Public Admin.
AI4EDU Institute Staff
MiXiaoTu (C2), iPad (C4)
[…]
[0.5pt/0.5pt] P14
Mother
3-yo daughter, 7-yo son
Comp. Sci.
Company Employee
HuoHuoTu (C2), Alpha Egg (C5), Tmall Genie (C6), XiaoAi (C6)
[…]
[0.5pt/0.5pt] P15
Father
3-yo daughter, 10-yo son
Psychotechnics
AI4EDU Institute Researcher
Interactive Talking Pen (C1), Storyplayer (C2), Alpha Egg (C5), Tmall Genie (C6)
[…]
[0.5pt/0.5pt] P16
Mother
4-yo son
Lawyer
Company Employee
Interactive Talking Pen (C1), Qiaohu, Tmall Genie (C6)
[…]
[0.5pt/0.5pt] P17
Father
3-yo daughter, 7-yo daughter
Education
Kindergartener
Tmall Genie (C6)
Table 1. Demographic information of our participants and the tools they have used. (C1-C6 used in the last column corresponded to the categories to which these technologies belonged, with details shown in Table 2.)
Of the 17 interviewees, 10 were mothers and 7 were fathers. 11 of them had one child between the ages of 3 and 6, and 6 had two children, with at least one falling in the age range of 3 to 6. Their educational and work backgrounds were varied, ranging from information technology, computer science, education, finance, law, and administrative management, to health. All of them had used at least one kind of AI-based intelligent tool for storytelling in their opinions. Table 1 shows demographic information of our participants and the tools they have used.
3.2. Data Collection
The interviews were semi-structured, including questions about 1) their daily storytelling practices and experiences, 2) the AI-based intelligent storytelling tools they have used and how they experienced these tools (when the technologies and tools reported by participants did not align well with research communities’ universal definition of AI technologies, we explored as much as possible the ”intelligent functions” that parents considered from these technologies and tools), 3) the benefits and challenges they experienced, and 4) their expectations of future intelligent tools. For parents who have used LLM-based tools (e.g. ChatGPT), we specifically asked 5) how they considered if LLM-based technologies were used to support preschooler-oriented storytelling. All interview questions were designed to be general so that the questions could be inclusive to participants to share various levels of experiences with different technology-mediated storytelling practices.
The interviews were mainly conducted in Mandarin by the first author. 12 were conducted via WeChat Voice call and Tencent Meeting, and 5 were conducted in person. Each interview lasted for 40-60 minutes, with a 100 CNY honorarium. During the interview, we encouraged participants to provide detailed demonstrations of how these tools functioned and how they used them. When some interesting points or prior experiences were mentioned, we asked appropriate follow-up questions for more details and concrete examples. With participants’ permission, we also obtained other forms of data to triangulate their interview responses. These data included 1) screenshots of mobile applications the participants have used and 2) related photos or videos they shot, if any.
3.3. Data Analysis
With the participants’ permission, we recorded the audio of all interviews and transcribed them into Chinese verbatim after the interviews for analysis. We followed thematic analysis as outlined by Clarke et al. (2015) to analyze the main themes in the interview data in an inductive way (Thomas, 2006). Three authors participated in the data analysis process. We started coding and analysis while the data was being collected. We iterated through generating codes and themes from collected data and checking and elaborating these codes by collecting more data. The data collection phase ended when recurring themes were identified.
During the open coding phase, three authors independently reviewed the collected data and generated codes pertaining to our research questions. We had regular meetings on a weekly basis to discuss codes and ensure reliability. After this stage, we generated the initial code list, capturing participants’ practices, motivations, and perceptions related to their daily storytelling actions; their usage, motivations, and experiences related to the intelligent storytelling tools they had used, as well as their perceptions and expectations to the general AI-based storytelling tools.
Based on this code list, we then re-focused our analysis at the broader level of themes, and synthesized and grouped the codes to create higher-level overarching themes. The initial analysis centered on themes related to storytelling tools parents had used, parents’ adoption, usage and experiences of these tools, the difficulties and challenges they have encountered, their children’s adoption, usage and experiences of these tools, parents’ perceptions and concerns about LLM-based AI tools, and parents’ expectations of future AI technologies. As analysis proceeded, we found that, due to our open criteria to “what AI tools should be in our opinion and what kinds of AI tools participants had used”, some tools reported by our participants, such as Huohuotu, story player etc., were not aligned well with research communities’ typical definition of AI technologies [^2] https://en.wikipedia.org/wiki/Artificial_intelligence on the one hand, and some, such as part of apps in iPad, were not used for story reading and telling, but for learning on the other hand. Considering the research aim of “gaining insights into parents’ practical adoption, use, experiences, and perceptions toward AI-based storytelling technologies and tools, through which to shed light on how we can better design future AI tools to support parent-preschooler interactive storytelling”, we removed codes regarding parents’ practical usage and experience of these not so AI-based tools and not storytelling-oriented tools to focus the analysis more on the content relevant to our research questions. In addition, we realized that the perceptions of using existing intelligent storytelling tools and LLM-based AI tools were closely related. We then further focused on themes that reflect similarities and differences between existing intelligent storytelling tools and LLM-based AI tools.
Through several rounds of analysis and discussions, three authors reached an agreement and generated our final satisfactory thematic map of the data, which included four primary themes: 1) parents’ attitudes and perception to AI-based storytelling with preschoolers, 2) their perceived benefits, and encounter 3) interactive and 4) algorithmic challenges in practice. In the following section, we present the full details of these themes. Throughout, we use representative quotes from our participants, translated from Chinese into English by the first author, to illustrate our points. To protect our participants’ identities, we use P1, P2, etc., to denote our study participants.
3.4. Positionality of the Authors
We acknowledge that, for qualitative research, the research perspectives and approaches are shaped by our own experiences and positionality (Pascoe, 2022). Specifically, we are researchers living and working in the U.S. and China, primarily researching human-centered AI and the perceived trustworthiness, fairness and effectiveness of AI in education. We come from a mix of disciplinary backgrounds, including Computer Science, Learning Sciences and Technology, Education, and HCI, which we have drawn on to conduct prior research into sociotechnical approaches to human-centered AI design practices. We have experience working with school teachers and district personnel on technology integration, and AI practitioners on projects related to human-centered educational technology design. In general, we recognize the tremendous potential of the rapidly developed AI in promoting societal progress and enhancing people’s quality of life and work. However, we do not agree with the notion that AI will replace human roles, such as the role of parents in children’s storytelling. We firmly believe that the optimal design approach should be human-centered, offering greater flexibility and adaptability. Our results confirm the validity of our hypothesis.
3.5. Ethical Considerations
All researchers received formal ethics training, and were granted official certificates from the ethics committee in their institute. During the study, we took careful steps to protect user privacy. Before we started interviews, we informed participants of our intention and background information, and got permission from them. All data collected during our study was used in an anonymized way, i.e., there was no link between the collected data and an individual user. We strictly abide by the confidentiality agreement, which promises the dataset is only used for this research and any form of information leakage is not permitted. Our institute does not fall under the jurisdiction of an internal review board or similar ethics committee. We did, however, get feedback from peers inside and outside our faculty to validate the ethical perspective of our research.
4. Findings
Table 2. Some AI-based storytelling technologies reported by our participants
Our findings generally suggested that it had been very common among parents to utilize various types of AI-based technologies and tools to facilitate their daily parent-children storytelling activities (as shown in the right column of Table 1). Among 17 parents we interviewed, 14 had used more than two types of technologies to support their daily parent-children storytelling practices. According to the interaction modes and functionalities, we roughly classified these tools into six categories, using C1-C6 to present. Specific details can be found in Table 2. Some of these tools were aligned well with research communities’ typical definition of AI technologies, while some others, such as tools of C1 and C2, were not. In the following, we will elaborate on the experiences, perceived benefits, and encountered challenges of our participants in using these tools.
4.1. Parents’ Attitudes and Expectations to AI-Based Storytelling Technologies
From parents’ perspective, our participants expressed mixed (positive and negative) attitudes towards existing AI-based storytelling technology. We found these attitudes were somewhat correlated with parents’ educational background, work experience, and parenting philosophies. Regardless of their attitudes, they all believed AI’s potential to support storytelling scenarios for preschoolers. However, they emphasized the functionalities it possessed and the role it played required careful attention. All of our participants emphasized that no matter how intelligent AI technologies become, they should be parents’ assistants in storytelling practices, rather than replacing parents to work. In the following sections, we will elaborate on parents’ attitudes and expectations regarding AI-based storytelling tools.
4.1.1. Parents’ Attitudes to AI-based Storytelling for Preschoolers
Our participants’ attitudes towards existing AI-based storytelling technologies varied from individual to individual. Some were generally positive and while others were relatively negative. Further, we found parents’ individual educational or occupational backgrounds influenced their individualized adoption and attitudes to AI-based storytelling tools to some extent. Among our interviewees, five participants were engaged in technology-related professions, or with technology-related educational background (P5 was a CIO of a company, P6 worked in the field of digital healthcare, P8 and P11 were involved in computational science research, and P13 held a managerial position in an intelligent education research institution). In contrast to other participants, these five interviewees exhibited more positive attitudes and higher expectations for future AI-based storytelling technologies. They believed that despite the current limitations of AI-based storytelling tools, there was great potential for future AI technologies to address these limitations and better assist parent-children storytelling. They expressed their eager anticipation to the emergence of such technologies. In contrast, four participants with educational backgrounds or occupations in the educational field (P2, P7, P12 and P17), hold relatively conservative attitudes toward AI-based storytelling tools. During the interviews with them, they explicitly expressed strong dislike and even disapproval of preschoolers’ use of AI-based tools, with the reasons of challenge between preschoolers’ limited discriminatory abilities and almost uncontrollable AI (P2), challenge between preschoolers’ cognitive development level and AI’s virtual interaction mode (P7), challenge between preschooler’ free, personalized development and AI’s overly direct goal orientation and standardization (P12) etc.
Meanwhile, parents’ parenting philosophies also significantly impact their attitudes toward accepting and using AI-based storytelling technologies. Throughout our interviews, participants often articulated their viewpoints using statements like “I believe that education for children should be XX; in this philosophy, AI XX.” Our data indicated that parenting philosophies had a more direct impact on parents’ acceptance and use of AI tools than educational or occupational backgrounds. For example, despite P2 having a background in computer science and P15 working in an intelligent education research institution, their attitudes toward AI tools are not particularly positive. Instead, they lean towards the idea of parents personally accompanying their children in storytelling, rather than relying on a smart machine.
4.1.2. Can AI Replace Parents’ Roles in Storytelling?
None of the participants believed AI technologies could replace the role of parents in storytelling scenarios because “it cannot support interactions with emotion and warmth, and cannot replace parents’ love, care and hugs” [P13, a father with a 3-year-old son and a 5-year-old daughter]. They emphasized that story reading and telling with preschoolers was a very children-centered and comprehensive process. Through the storytelling process, they hoped children could perceive the world, learn different knowledge and skills, and feel different cultural contexts. At the same time, they also pointed out that it was not only for learning and education, but also for guiding children’s active thinking, cultivating their reading habits, or just be there and spend parent-child time with my child [P5].
Yet, AI-based storytelling tools paid more attention to reading the story or teaching knowledge. They couldn’t be able to care for and love them as their parents did. Even the more powerful LLM-based applications couldn’t replace parents to carry out such a children-centered comprehensive process. P9, a mother with a 6-year-old son strongly opposed the existing trend of using AI technology to tell stories for preschoolers. She believed that children’s growth should be accompanied by parents, rather than a smart machine.
“You propose to design a super smart machine and ask me to use it to educate my son. Why? I don’t understand, I think the most important thing is the interaction between parents and children, between human and human. Mothers should not find a machine to accompany their children, but rather ‘I accompany you’. A machine is just a machine, no matter how advanced it is, it is still just a machine.”[P9]
At the same time, many of our participants thought that AI-mediated storytelling was not real storytelling or reading. They emphasized that storytelling and reading in the real world was not just about telling and reading those stories, asking the questions, or teaching some knowledge, but also the reading habits, thinking methods, and even postures and attitudes of reading. Parents need to pay attention to all of these aspects, but AI tools often only read the stories.
While all our participants agreed that AI technologies absolutely couldn’t replace the role of parents in practical storytelling scenarios, most of them explicitly expressed their expectations of intelligent technologies to become their “assistants” or “secret weapons”, with functions of freeing up parents’ physical labor (e.g. stroybook reading), helping them generate valuable and high-quality interactive content or questions, or inspiring them on how to interact with their children.
In summary, our participants expressed varied attitudes toward AI-based storytelling but shared similar expectations of it being parents’ assistant. Then, do the existing AI-based storytelling tools meet parents’ expectations? How effective are they in practical use? What is the parents’ experience with their usage? We now discuss these questions in more detail in the following sections.
4.2. Perceived Value of Using AI-Based Storytelling Technologies
When reflecting on the experience of using AI-based storytelling tools to associate daily parent-child storytelling, our participants talked about some benefits of these technologies they perceived, which will be presented in this section.
To be specific, most of them expressed that they completely agreed that interactive storytelling was a very important and beneficial parent-child activity for preschoolers. Yet, for most of them, this conversation-rich interactive storytelling with preschoolers was full of practical challenges. For example, P9, a mother with a 6-year-old son, said, “I really want to interact with my son through expanding the story content, but I don’t know what content should be expanded. I even can’t understand the questions he asks me.” Some participants also noted that engaging in daily, high-quality interactive storytelling, as they expected, represented significant challenges in terms of time, energy, and physical stamina. This was particularly true for today’s young parents who were often immersed in fast-paced and demanding work.
Under such a situation, AI-based storytelling tools, worked as more captivating and technologically advanced toys, could help alleviate these challenges to some extent. Our participants expressed that although existing AI-based technologies they have used didn’t fully meet their expectations in supporting children-oriented interactive storytelling (see more details in the following subsections), they did find them helpful in interacting with children when they were occupied (i.e., playing with children or telling stories). Meanwhile, these technologies could partially relieve parents from those repetitive and monotonous story-reading tasks, allowing them more time to engage in the desired high-quality interactions. As P6, a mother of a 6-year-old daughter, said,
I wish I could spend an hour every day telling her stories, but I don’t have enough time and energy. I can’t always be present with my child. So, how can my child spend that time? That’s where XiaoAi (a smart speaker similar to C6 in Table 2) comes in. At least it can read stories to her, or engage in conversations. Although it doesn’t work very well at the moment, it allows me to dedicate my efforts to more challenging tasks, such as interacting with my child and responding to her questions. ” [P6]
At the same time, many of our participants stated that compared to plush toys and the like, these AI-based tools were relatively “better” toys for children, because they could bring “novelties and a sense of technology” [P8, P13] to children, which parents couldn’t bring. Most existing storytelling tools had some children-tailored interactive designs, such as immersive storytelling processes that blend multiple modalities (“Children could touch, hear, see, and talk with it.” [P6]), and vivid presentations (“The sound of the storyteller is very pleasant and full of emotions and varied intonations, and the animals can move and make sounds” [P1]), which worked well to attract children to participate in the interaction, improving their storytelling experience significantly.
In addition, most of these technologies integrated more or less learning resources, or had some targeted interactive designs, which allowed children to learn something or develop certain abilities while playing. As P8, a mother of 6-year-old son, said: “You always need to buy something for the child, right? Do you buy a bunch of Ultraman or cloth dolls, or these intelligent things? No matter how the practical effect is, it provides the possible opportunities to cultivate the creative thinking abilities of children, or learn something when playing.” [P8]
Despite these perceived benefits, our participants considered that most of the storytelling tools they had bought and used were hard to consistently engage preschoolers and often struggled to achieve the intelligent features claimed by the products. The practical effectiveness and efficiency in supporting quality storytelling were limited. Our participants reported several challenges with these tools that hindered their practical effectiveness. We generally categorized these challenges into two types - interactive and algorithmic, and detailed them in the following two sections.
4.3. Interactive Challenges in Supporting Storytelling Aligned with Cognitive Characteristics
The first category of challenges we identified was interactive challenges, mainly referring to challenges and dilemmas caused by interactive ways and interfaces existing AI-based storytelling tools used, which primarily arose from existing tools’ insufficient consideration of preschool-age children’s cognitive characteristics and their limited abilities to accept and use these technologies.
4.3.1. Preschoolers are Difficult to Use AI Tools Proactively, Independently, and Consistently
According to our participants, preschoolers often exhibited short attention spans, lacked self-learning and self-management abilities, etc. These characteristics made it hard for them to use various AI-based tools actively, independently, and consistently. Most of the participants thought it was unrealistic to expect preschoolers to use these tools to read a storybook independently. Instead, parents’ engagement was crucial in determining the practical effectiveness of these tools.
To be specific, most existing AI-based storytelling tools require specific usage processes and scenarios. For example, the use of the interactive talking pen (C1 in Table 2) required parents to “take out the pen, turn it on, find the corresponding book, and flip to the specific story [P8]” and then the children could use it to read a story. “Without parents’ operations and guidance, preschoolers are unlikely to proactively use these tools to read the story” [P8]. As for the interactive and question-answering function, our participants expressed that when their children had questions or hoped to interact, asking and interacting with parents was an easier and more straightforward way for them.
Further, preschoolers’ short attention spans made it challenging to consistently and sustainably use any tools with focused attention. Consequently, whether and how parents were involved directly affected the practical effectiveness of tools. Many parents in our study expressed that there was a notable difference in effectiveness between children using these tools independently and when accompanied by parents. P5, a mother of a 6-year-old daughter, considered that, “if parents are not involved, even for the best product, it is difficult to sustain children’s interest and attention, and motivate them to read a story with it for 15-20 minutes steadily”. P8, who had purchased many AI-based storytelling tools for her son, shared her experience with these tools.
Our (parents’) involvement or not has a completely different using effect. I can let these tools tell stories to my son. Tools, like Luca (C5 in Table 2) and Xivo (similar to C4 in Table 2), are well-designed for this task. Objectively, they perform well in terms of design, interaction, and storytelling. The issue is, preschoolers have short attention spans, without any active learning consciousness. It is their basic characteristic. … I’ve heard that some children start to slack off once they’ve mastered the tool. They can quickly complete the reading task through repeatedly clicking the progress bar.” [P8]
4.3.2. Preschoolers Need the Sense of Authenticity and Constructiveness
The virtual interactive modality, without the sense of authenticity and constructiveness, was also reported by our participants as one major reason that current AI-based storytelling technologies couldn’t sustainably attract children’s attention and achieve the expected storytelling effect. They expressed that preschoolers were still in the stage of development of motor thinking, and their abstract thinking ability was not well developed. This thinking development stage determined that they needed more physical and real interactive processes with a sense of constructiveness, instead of the virtual way, as most existing intelligent tools used. They expressed that the virtual interactive way brought certain cognitive difficulties for preschoolers to form cognition in their minds. P15 used his son’s experience of playing with building blocks to explain this sense of authenticity and constructiveness:
My son can play with building blocks for more than two hours. He has a sense of reality that he is really building something. Whatever he thinks in his mind, he can immediately construct it, such as a castle, a kitchen, or a gun [P7]”.
Meanwhile, compared to human-human interaction, our participants also recognized that virtual interaction methods lacked the necessary emotional connection, which resulted in intelligent agents struggling to effectively sustain children’s attention and engagement. As P15 said,
I purchased two expensive robots. In the beginning, there was a sense of novelty. My child interacted with them actively, saying things like ‘Egg, come with me; Egg, sing me a song; Egg, tell me a story.’ But he only played with them for three days. Currently, they are just expensive toys, sitting in the corner and serving no purpose. Children can’t establish an emotional connection with them. I think things that can be touched and felt will provide more satisfying experiences for children. [P15]
Additionally, our participants also considered systematic and contextual information as significant in constructing children’s cognition of the story, but lacked in AI-based storytelling. Consequently, it became challenging to use AI-based storytelling tools to support attractive storytelling for children. They stated that in traditional paper books-based storytelling, the storybooks often possessed coherent story structures and contextual situations. Throughout the reading process, children could perceive this contextual information and construct the stories in their minds through flipping the pages back and forth. In contrast, AI agents only presented the story plot at a specific moment, devoid of any contextual information. This limited their ability to develop children’s comprehensive understanding of the story, subsequently impacting their engagement and interests.
4.3.3. Preschoolers Need Interaction, Instead of a Correct Answer
Interaction and stimulating children’s active thinking were reported by our participants as a primary expectation of storytelling, even more important than learning knowledge. Through storytelling, they desired children would be guided to actively participate in the process of imagination, emotional expression, and critical thinking, which were helpful to enhance their creativity, empathy, and problem-solving skills. P13 showed us a vivid storytelling process he conducted with his 5-year-old daughter, which illustrated this interacting and inspiring purpose very well.
One day, I told my daughter the story of ‘monkey move corn’, in which the monkey initially held the corn. After seeing a peach, the monkey held the peach and threw the corn away. After seeing a watermelon, it then held the watermelon and threw the peach away. … In the end, the monkey didn’t catch anything. This classic story has many inherent truths, such as the importance of having a clear goal and being focused on it. However, I didn’t tell my daughter this moral directly. Instead, I asked her, ‘if you were the monkey, what would you do?’ To my surprise, my daughter responded with great insight. She said, ‘if I were the monkey, I might have three ways to solve the problem. First, I would bring a bigger bag to take everything. Second, I would pick corn today, peaches tomorrow, and watermelons the day after tomorrow. Third, I would call my friends to come with me, then we could gather corn, peaches, and watermelons together.’ [P13] ”
Based on this story, P13 and many other participants emphasized that meaningful interactive storytelling should aim to stimulate children’s active thinking, rather than simply imparting a fixed moral or lesson. However, most current AI-based storytelling tools pay primary attention to supporting correct question-answering, i.e., providing a ‘correct answer’ to children’s questions. They were completely unable to support their desired interactive storytelling. P9, for instance, described her perception of using Tmall Genie (a general conversational agent, similar to C6 in Table 2) to support storytelling practices for preschoolers:
When the child asks why the sky is blue, it gives an answer and then ends the conversation. This is of limited meaning to the child. It cannot inspire children’s interests and thinking. They may not understand it, and then their interest stops there. What preschoolers need is actually an interactive process, not just a correct answer.” [P9]
For this reason, some of our participants strongly rejected the tools with explicit educational purposes, because they might limit children’s thinking and self-reflection.
Pre-school storytelling is not mainly for learning knowledge, but many existing tools focus on ‘teaching you something’. They are very eager to impart knowledge to children. I don’t like them at all. A story is just a story. You just need to tell a story. What children can learn or think from the story depends on the interactor and themselves. ”[P12]
Taken together, our participants illustrated a series of interactive challenges between the interactive ways existing AI-based storytelling tools are used and the interactions aligning with preschoolers’s cognitive and behavioral characteristics. They suggested that AI-based storytelling technologies should carefully consider preschoolers’s inherent characteristics and cognitive development level. Meanwhile, technologies should guide children’s active thinking, instead of telling them the correct answer or specific knowledge directly.
4.4. Algorithmic Challenges in Supporting Child-Centered, Intelligent, Personalized, and Heuristic Storytelling
The second category of challenges we identified was algorithmic challenges, mainly referring to challenges and dilemmas caused by AI algorithms embedded in storytelling tools. These challenges primarily arose from the limited AI algorithms in supporting child-centered, heuristic, personalized, and adaptive storytelling practices. As a result, the usability and effectiveness of these tools in facilitating practical, parent-expected intelligent storytelling for preschoolers were constrained.
4.4.1. Challenges in Supporting ‘Intelligent’ Interaction
Silly and limited-intelligent interaction was one major algorithmic challenge our participants experienced when they used AI-based tools, no matter children-tailored tools (e.g. C3, C4, C5 in Table 2) or general intelligent tools (e.g C6 in Table 2), to support the storytelling with their children. For children-tailored tools (e.g C1-C5 in Table 2), for instance, our participants commonly considered their intelligence was limited, because their interaction ways and content were notably inflexible, adhering to a fixed structure and only permitting content access and updates within the limitations set by the merchant. The experience of P1 using Little Genius (similar to C4 in Table 2) illustrated this limited intelligence very well: “its content is almost static, and the interaction remains unchanged. Whether you open it this year or next year, it remains the same. It lacks adaptability and fails to adjust according to the child’s developmental needs.
For general intelligent conversational agents based on massive corpora, such as C6 in Table 2, our participants considered that, although they performed commands well, their abilities in understanding children’s attention and supporting children-oriented dialogue and interaction were still very limited. P6’s experiences of using XiaoDu (C6 in Table 2) illustrated this limited intelligence very well:
Whether it’s XiaoDu or XiaoAi, its intelligence level is not that high. It often struggles to understand a child’s command and falls short in delivering appropriate responses, especially when the child’s speech is unclear. … For instance, when you ask it to narrate a story about Journey to the West, it might veer off-topic and present various unrelated products related to Journey to the West.” [P6]
In particular, our participants highlighted that in real-life scenarios, children’s interests and questions were frequently diverse. However, current AI-based storytelling technologies are more like search engines. They could effectively answer questions that could be queried through search engines, but failed to answer those that couldn’t be queried. For example, P8 evaluated the use of Alpha Egg (C5 in Table 2):
The function is not particularly smart. It’s just a search engine. When you ask who is Osa, it searches online and reads the answer to you. Once your question can’t be searched online, then it fails to answer.” [P8]
Even with more intelligent generative AI technologies, parents in our study still expressed their concerns about their capacity to genuinely grasp and decipher preschoolers’s intentions and furnish responses. They expressed uncertainty about whether and how generative AI technologies could answer their children’s endless and personalized ”why” questions well, because many of these questions actually didn’t have standard answers.
Child’s questions are often very strange. They may ask why witches are female and not male, why witches have pointy noses, and why witches are thin instead of fat. Or ask why Harry Potter is named Harry Potter, why he rides a broomstick instead of a vacuum cleaner. There are no standard answers to these questions. How can machines answer these questions? If the machine asks the child, ‘do you think it’s right or wrong for the Yellow Robe Monster to steal someone’s wife and treat her especially well?’ and the child answers with their reasoning, how should the machine respond?” [P6]
4.4.2. Challenges in Supporting Children-Appropriated Interaction and Content Generation
Whether the AI tools could support children-appropriated interaction or generate children-appropriated content was another significant algorithm challenge reported by our participants. Our participants indicated that when they used tools such as C5 and C6 in Table 2 to answer children’s story-related questions, the generated answers were often serious and professional, not specifically tailored to children’s cognitive level and psychological age, which caused it difficult for Children to understand. P15’s experience of using XiaoAi to answer children’s questions illustrated this very well:
My child asked me what ‘shame’ was in the story, and I asked XiaoAi. It gave an explanation. But it was like a dictionary, just reading out the explanation in the dictionary. If you ask it what a water tank is, it can also give a very scientific answer. But such kind of answers were too serious for the young child to understand.” [P15]
A similar concern was also reported by participants who had used ChatGPT to answer children’s questions. They expressed that the answer generated by ChatGPT was not very suitable for preschoolers, particularly in terms of knowledge level, language structure, and presentation. As P8 said, “The content is often too long and professional. Many terms are unfamiliar to preschoolers. If it is used to talk to children, it may need to be shorter and simpler. I think this is a technological challenge. There are thousands of questions, and how can it answer them in the concise and appropriate language that children can understand?
4.4.3. Challenges in Supporting Personalized Interaction
Personalized interaction was reported by our participants as the basic interaction requirement of AI-based storytelling technologies. They emphasized that, in practice, each child’s characteristics, interests and questions were very different, “Some children are receptive, while others are more proactive” [P10], and “some children prefer to read by themselves, while others like you (the parents) to tell them” [P5]. These personalized features caused that, even if children saw the same story, their points of interest might be completely different. As P15 said, “with the same story, different children may focus on different things. Boys and girls think differently, and my child and her tablemate in kindergarten also think differently”.
In addition to differences among children, our participants emphasized that one child’s character, interests and focuses also dynamically change. Many participants told us that their children read the same story repeatedly, but their understanding and points of interest about the story content changed continuously as they grew older. However, although many existing AI-based storytelling technologies claimed that they supported personalized interaction and learning, in our participants’ view, current personalized interaction was very coarse-grained (e.g., “first graders may read these books, second graders may read these books” [P12]), which couldn’t fully support their expected personalized interaction with preschoolers. They emphasized that true personalization for preschoolers needed to consider various factors such as children’s real-time interests, needs, characteristics, personalities, etc., not just the grade and age. Meanwhile, these technologies didn’t provide the functions of automatically perceiving personalized characteristics of children. Instead, they often relied on parents manually inputting some basic characteristics of the children, such as age, gender, interest, etc., and based on these characteristics to conduct “coarse-grained personalization”. This made it more difficult for the so-called personalized features to become a reality in practical storytelling scenarios.
Some technologies employed recommendation algorithms to support personalized story and content recommendations. Yet, our participants considered that the current recommendation algorithms were almost incapable of providing human-centered personalized recommendations. They worked more like “the child listens to a story today, then the system will recommend the similar stories for the next few days” [P4]. P4 considered these recommendation algorithms as “greatly limiting the personalized development of preschoolers”. Our participants agreed that perceiving children’s personalized characteristics and interacting with children based on their personalized characteristics were highly challenging, even for very intelligent LLM-based tools such as ChatGPT. “Large models might know the overall patterns, but didn’t know the specific feature and need of each child”, as P7 said.
4.4.4. Challenges in Supporting Situated and Adaptive Interaction
According to our participants, the practical storytelling was a dynamic and situated process. During the practical storytelling processes, parents frequently had a dynamic awareness of children’s real-time interests and attention levels, based on what to dynamically adjust the interaction contents. Furthermore, our participants expressed that even as parents, they may not have a comprehensive and correct understanding of their children’s characteristics and interests. Their understanding of their children and the interactive ways and contents they chose were gradually discovered and adjusted during the interactive process.
We also didn’t know what our child liked at the beginning. We can only adjust the interaction based on his feedback and state. For example, from his eyes and body language, we can learn that he doesn’t want to read anymore, he wants to go out to play, or he is not interested in this story. If we ask a question and he doesn’t react, we will change the question or the wording. Through this process, we gradually know what content and interactive methods he likes.” [P9]
Nonetheless, our participants stated that the majority of existing AI-based storytelling technologies were unable to dynamically perceive children’s real-time situations very well, which impacted the actual effectiveness of these tools directly. “Machines have a lot of knowledge and just want to impart it to children. They have no idea whether children want to listen or are interested in what they are saying.”, as P12 said.
4.4.5. Challenges in Supporting Children’s Active and Diversified Thinking
Throughout our study, our participants consistently emphasized that every child was unique. However, current technologies, they considered, fell short of adequately supporting the diverse development of children. For children-tailored conversational agents (e.g. C4 and C5 in Table 2), which were built on specialized children’s corpora and could ask children questions proactively, the interaction was primarily based on relatively fixed patterns and language models, which evaluated by our participants as “limiting children’s active and diversified development and thinking” [P11]. They thought children should engage actively in thinking and posing questions based on their curiosity, rather than relying on a machine to ask pre-determined questions and provide predefined answers. P4 explained this opinion through his using experience to C4 in Table 2:
Each story with it is followed by a universal truth, such as the stepmother is evil. But the child may be interested in stepmother’s beautiful dress, and its color. The process of proactive thinking and questioning is valuable. Even though many of the child’s questions may appear strange, silly, or stupid, they are actually their own questions and expressions of their interests.” [P4]
For general intelligent conversational agents based on massive corpora, such as C6 in Table 2, and even more intelligent generative AI technologies such as ChatGPT, many of our participants, such as P4 and P7, still expressed their rejection of using it to support children’s storytelling and interaction, with concerns of “standard answers” given by ChatGPT limiting the diverse development of children. They thought that, no matter how intelligent, AI technologies were fundamentally built upon extensive data-driven pattern recognition, i.e. mining rules based on massive data and providing ‘standardized’ and ‘correct’ answers, which were definitely not suitable for preschoolers, and even potentially mold them into fixed paradigms.
I don’t support using ChatGPT to tell stories or interact with children. ChatGPT can answer many questions, hundreds of thousands of questions, but these questions and their answers have already been summarized by predecessors. It standardizes these things and gives them to children, which actually limits children’s diversified thinking and development.
4.4.6. Challenges in Supporting Authenticity, Civilized, Safe and Ethical Interaction
Challenges in supporting authenticity, civilized, safe, and ethical interaction primarily pertained to general AI technologies such as C6 in Table 2, and in particular, the booming generative AI technologies due to their lack of transparency and insufficient controllability of LLM-based technologies (Wu et al., 2022). Our participants expressed that preschoolers did not have the ability to discern multiple information. Given the insufficient controllability, they were worried about what kind of conversations children would have with such a smart robot, and who could ensure the authenticity of the output content.
The value orientation of the content output by ChatGPT was also another concern expressed by our participants. P2 was a Christian and she explained this ”value orientation” to us: “In our family, we prioritize this aspect as we have strong beliefs. We focus on concepts that align with our faith, and scientific knowledge that contradicts our beliefs is not widely accepted. For example, we adhere to the creation theory rather than the evolution theory. I am unsure whether ChatGPT has the capability to accommodate or respect these specific preferences”[P2].
In addition, parents in our study were also concerned about LLM-based AI technologies’ ability to recognize and guide children correctly when they asked inappropriate questions. As P2 said, “if a child asks how to not do homework or how to beat my classmate, how would the machine respond, and can it allow the children to ask that question?” [P2]. The civility of language was also a concern for parents. Given children’s strong learning and imitation abilities, our participants worried if ChatGPT used uncivilized or inappropriate language styles, it would have a very negative impact on children. Based on these concerns, many of our parents stated that they currently did not allow their children to access uncertain content on the Internet without their control. “Nowadays, many things seem uncontrollable, and the world is very dangerous. I will never let my children face the evil world without my control” [P2].
5. Discussion
This paper has illustrated how parents with preschoolers use, experience, and perceive the existing AI-based storytelling technologies and tools in practice. Our findings indicated that, while parents perceived some value when using these tools, they expressed that existing AI-based storytelling tools they had bought faced numerous practical challenges in supporting their expected storytelling. We categorize these challenges into interactive and algorithmic and elaborate on them in the Finding section. According to these findings, we now delve into parents’ expectations for ideal parent-child interactive storytelling and propose research implications for future AI-based storytelling research.
5.1. Re-Conceptualizing the Interaction in AI-Mediated Storytelling Scenarios: Taking Parents as “Beneficiary”
Most existing AI-based storytelling and related systems primarily focus on children-AI interaction. However, our research has revealed several challenges with this children-AI interaction paradigm. Specifically, preschoolers typically lack the suitability to use tools for storytelling independently and consistently, making solo storybook reading and telling with AI a challenging task. Meanwhile, the capabilities of existing AI tools are insufficient to meet parents’ expected storytelling. Furthermore, the process of storytelling for preschoolers is inherently a parent-children interactive experience, with one primary goal of enhancing parent-child relationships, as reported by our participants as well as existing literature (Wright, 1995; Chang and Huang, 2016; Peck, 1989). However, this aspect is not well-supported under existing Children-AI interaction paradigm.
In CSCW, researchers are increasingly acknowledging the limitations of this Children-AI interaction paradigm (Zhang et al., 2022; Harbig et al., 2011), and also emphasizing the positive impact of parental involvement in children’s use of AI for storytelling (Cagiltay et al., 2022; Zhao and McEwen, 2021). Our study, echoing this literature, further emphasizes the importance of recognizing parents as direct beneficiaries of AI tools, and supporting an interactive paradigm among parents, children, and AI tools, rather than focusing solely on children-AI interaction. Based on our findings, We specifically propose the following design implications.
Taking Parents as Beneficiaries: Support, not Substitute. We argue that the role of AI tools should be to assist parents in storytelling, rather than substituting their role entirely. According to our participants, despite the limitations of existing AI tools, they expressed expectations for AI tools to assist them by, for instance, alleviating some of the physical labor involved (e.g., helping them read stories), generating valuable and high-quality interactive content or questions for them when they lack expertise, providing inspiration for interactions with their children, and preparing appropriate resource for supporting their storytelling with children. These supporting functions could be considered by future AI tools.
Taking Parents as Beneficiaries: Beyond a Storytelling Toolman. By taking parents as beneficiaries, we also suggest shifting parents’ role in AI-based storytelling from merely tool man or technology supporters, as most existing tools do, to active participants. Activities such as role-playing and interactive play can be incorporated. While role-playing is already utilized in some tools (Zhao and McEwen, 2021; Dietz et al., 2021), many of them still take parents as supporters in terms of helping with media content selection and text-to-speech usage, rather than fully participating as players. We recommend treating parents as true players to enhance their involvement. Ongoing research on interactive technologies (e.g., (Chen et al., 2022)) may offer potential support for better parent-child interactive storytelling experiences.
Enabling Parents’ Personal Experience Through Co-Creating.. According to most of our participants, during their actual storytelling with preschoolers, their conversations often go beyond the storybooks, and integrate a lot of personal viewpoints and experiences, which has also been recognized in the literature as an essential and valuable way of storytelling (Vasalou et al., 2020; Zhang et al., 2022). We therefore propose AI-mediated storytelling could empower parents by involving them as co-storytellers, and providing them interfaces to seamlessly incorporate their personal experiences, childhood memories, or whatever they want into the narrative. Such integration could also effectively contribute to the realization of personalized story themes that align with parental objectives.
5.2. Design Considerations for Child-Appropriated Interaction: Enabling Preschoolers’ Psychology and Cognitive Features
In terms of interactive modality, our study has demonstrated that preschoolers often face certain cognitive difficulties in accepting the virtual interactive modality; Meanwhile, current AI tools often fail to effectively support preschoolers’ proactive and consistent participation. The main reason for these issues is that the current designs do not take into account the preschoolers’ cognitive and psychological features, resulting in a misalignment between child cognitive development and the virtual interactive way (Xu and Warschauer, 2020b).
According to these findings, we propose that future AI storytelling tools need to fit with preschoolers’ psychology and cognitive characteristics. Some classical theories of child cognitive development, such as Piaget’s theory of cognitive development (Piaget, 1976), could be used to guide the interactive design tailored for preschoolers. Within the framework of Piaget’s theory, each stage of development has distinct cognitive characteristics. Preschoolers are in the preoperational stage, where their cognitive processes remain anchored in concrete experiences, starting to be symbolic thinking. They can use imagination and symbols to express and solve problems, but their grasp of logical reasoning and abstract concepts is still limited(Piaget, 1976).
Given the cognitive and psychological feathers of preschoolers presented by Piaget (1976), as well as our findings, we propose concrete, multimodal, and tangible interaction modalities for future AI-based storytelling tools for preschoolers, such as providing tangible user interfaces (e.g., physical objects) for virtual interactions, incorporating motion perception, to enhance user engagement; Supporting storytelling processes with concrete experiences (e.g. AI building blocks (Cai, 2023b)) to capture children’s attention; Enabling preschoolers to engage in storytelling through a combination of real and virtual interactions (e.g. creating 3D story environments where children can manipulate virtual characters physically to accomplish story narrations (Wang et al., 2014)). These approaches provide a foundation for concrete thinking in children at the preoperational stage and, with the aid of virtual interaction modalities, facilitate the development of symbolic thinking.
It is noteworthy that we are not denying the value of virtual interaction modalities. Instead, similar to existing studies (Xu et al., 2021b; Pantoja et al., 2019), our participants also reported benefits of virtual (e.g. voice-based) interaction in enhancing children’s language abilities. What we want to emphasize here is the importance of combining multiple interaction modalities. That is, integrating virtual interaction modality into tangible interfaces, serving as a medium for children to transition between concrete thinking and symbolic thinking.
5.3. Design Considerations for Child-Appropriated AI Algorithms: Carefully Considering ”Appropriated”
In terms of core technologies and algorithms of existing AI-based storytelling tools, our participants reported significant challenges, primarily regarding the silly, unintelligent algorithms, and their failure to output adaptive and understandable content to children and answer their ”a hundred thousand whys”. From the technical perspective, a natural improving suggestion is to construct child-appropriated datasets and corpora that cater to preschoolers’ language and cognitive abilities, and then use them to enhance algorithm performance, making the output better suited to preschoolers’ needs in terms of language style, content, complexity, etc. Ongoing efforts have started to focus on these aspects, such as constructing question-answer storytelling datasets tailored to children’s needs (Xu et al., 2022; Chen et al., 2023), utilizing Knowledge Graph-based approaches for knowledge expansion (Chen et al., 2023), and improving AI to generate story book related, adaptive conversation (Yao et al., 2022; Kory and Breazeal, 2014). We believe that these research directions are crucial for the future AI-based storytelling technologies.
Moving forward, we further suggest carefully considering what is ”appropriate” for preschoolers and, based on that, exploring directions for algorithmic improvement. When we refer to ”children-appropriated,” we are considering whether the interaction and generated content are suitable for children. This entails aligning with their cognitive level, personal abilities, interests, preferences, and other relevant factors (Moyer et al., 1987). However, from the algorithmic perspective, the ”appropriate” is often reflected in the technical performance, e.g. the accuracy of answers, which the majority of our participants believe may not be applicable for preschoolers, because there is no ”standard or correct answer” for each child. For instance, algorithms may focus on whether it can generate the correct answer “poisoned apple” for the question “what did Snow White eat and die from?”, while children may be more interested in questions like “What dress was the witch wearing?” We therefore suggest future algorithms to support more open and diverse interactions that encourage children’s creative and critical thinking, rather than solely focusing on providing the ”accurate” answer. The advancing large language models have the potential to support such conversational needs, and researchers have also started to examine its potential in supporting children’s creative expression and storytelling (Han and Cai, 2023a, b). We believe it is a valuable direction, but need to carefully consider content safety and ethical issues which are significant concerns raised by our participants.
To conclude our discussion, we emphasize that not all challenges raised by parents can or should be resolved through AI algorithms, because, given AI’s understanding of ”appropriate”, it might be insufficient to answer the countless ”why” questions from preschoolers that have no correct answer, and support parents expected storytelling and personalized interactions. Further, AI solutions will inevitably involve the use of more perceptual technologies to gather additional data, posing challenges to the protection of children’s privacy. Our participation also expresses similar concerns in this regard. Echoing the discussion in section 5.1, we further emphasize it is crucial to carefully consider the role and functionality of AI tools in preschoolers’ storytelling - these tools should serve as aids for parents, enabling them to better support the storytelling process in a manner that aligns with their children’s unique needs.
5.4. Limitation and Future Work
Overall, our research presents an in-depth and empirical investigation of parents’ experiences and perceptions of AI-based storytelling. In particular, we elaborate on a range of challenges our participants reported in using existing AI-based tools to support their expected storytelling for preschoolers in practice. A major limitation of our study is the small sample size, which may result in possible sample selection bias (Heckman, 1979) and sample diversity (Wolda, 1981) issues. In particular, during the recruitment process, we did not consider the diverse characteristics of parents. However, during the research, we discovered that participants’ individual characteristics, such as their economic level, educational background, and professional expertise, significantly influenced their attitudes and perceptions toward using AI for storytelling. Unfortunately, given the sample size of our study, our current study did not analyze these perspectives very well. Furthermore, it is important to note that our study primarily took place in China. While this specific socio-technological context provided us with a unique opportunity to gain insights into AI-based storytelling practices, it also may limit the generalizability of our findings. In future research, we aim to expand our focus to a more diverse range of parental perspectives, painting a more complete picture of our research questions.
6. Conclusion
Using AI-based tools to support parent-children interactive storytelling has been a contemporary trend. In this work, we conducted an interview study to deeply examine how they use, experience, and perceive AI-based storytelling technologies in practical storytelling scenarios. Our findings suggest current AI-based storytelling technologies face numerous challenges in supporting storytelling for preschoolers as parents expect. We categorize them into interactive and algorithmic challenges, and elaborate on them in this paper. Drawing on these findings, we identified vital design implications to better facilitate the increasing demand for AI-based children-parent storytelling.
With the widespread applications of artificial intelligence in contemporary society, human-centered AI has become a focal point in CSCW, HCI and AI fields. However, how to effectively implement the human-centered vision and create practical technologies that truly put people first has consistently posed challenges. Our research focuses on the specific AI application scenario – AI-based storytelling for preschoolers, and delves into how the concept of human-centered AI can be realized in this scenario. While our research focuses on the specific scenario of storytelling, we believe that the design principles we discuss can be generalized to a broader range of technology designs aimed at facilitating children-AI interactions.
Acknowledgements.
To be added.